Two rectangles have the same width. The length of one is 1 foot longer than the width. The length of the other is 2 feet longer than the width. The larger rectangle has 5 more square feet than the smaller. What is the width of the rectangles?

Answer by ankor@dixie-net.com(22740)

Two rectangles have the same width.
The length of one is 1 foot longer than the width.
The length of the other is 2 feet longer than the width.
The larger rectangle has 5 more square feet than the smaller.
What is the width of the rectangles?
:
Let x = the width of the 1st & 2nd rectangles
then
(x+1) = the length of the 1st rect
and
(x+2) = the length of the 2nd rect
:
2nd area - 1st area = 5 sq/ft
x(x+2) - x(x+1) = 5
x^2 + 2x - x^2 - x = 5
combine like terms
x = 5 ft is the width of the rectangles
:
:
See if that checks out, find the area of each rect
2nd rect: 7 * 5 = 35
1st rect: 6 * 5 = 30
--------------------
Area difference: 5 sq/ft